can someone please show me how to do this problem:
Use the Gauss-Jordan method to solve the system of equations.
x+y+z=7
x-y+2z=7
2x+3z=14
a. 3z+14/2; z/2; z
b. -3z-14/2; z/2; z
c. -3z+ 14/2; z/2; z
d. -3z+14/2; 2z; z

Answer by stanbon(48568)

x+y+z=7
x-y+2z=7
2x+3z=14
---------------
Subtract the 1st row from the 2nd
Subtract 2 times the 1st from from the 3rd
-----------------------------
x + y + z = 7
0 - 2y + z = 0
0 - 2y + z = 0
- --------------
Subtract the 2nd row from the 3rd row to get:
x + y + z = 7
0 - 2y + z = 0
0 +- 0 + 0 = 0
- --------------
Conclusion:
z = 2y
---------
x + (1/2)z + z = 7
x = (-3/2)z + 7
----------------------
Solution:
x = (-3/2)z+7
y = (1/2)z
z = z
===========
Cheers,
Stan H.
=============
Ans: "c"
=============

a. 3z+14/2; z/2; z
b. -3z-14/2; z/2; z
c. -3z+ 14/2; z/2; z
d. -3z+14/2; 2z; z